Let L be the line given by < 3, −1, 2 > + t < 1, 1, −1 >, for t ∈ R.

(4.1) Show that the above line L lies on the plane −2x + 3y − 4z + 1 = 0.

(4.2) Find an equation for the plane through the point P = (3, −2, 4) that is perpendicular to the (4) line < −8, 2, 0 > + t < −3, 2, −7 >


1
Expert's answer
2021-07-06T09:19:58-0400

"<3,-1,2>+t<1,1,-1>\\
x=3+t, y=-1+t,\\
z=2-t"

(4.1)

To show that the Line lies on the plane −2x + 3y − 4z + 1 = 0

"-2(3+t)+3(-1+t)-4(2-t)+1\\
-6-2t-3+3t-8+4t+1=5t-16=0\\
t=\frac{16}{5}\\
x=3+t=\frac{31}{5}\\
y=-1+t=\frac{11}{5}\\
z=2-t=\frac{-6}{5}\\
-2x+3y-4z-1=-2\times \frac{31}{5}+3\times \frac{11}{5}-4\times \frac{-6}{5}\\+1=0"

Hence it lies on the plane -2x+3y-4z+1=0

(4.2)

"r_o=<3,-2,4>\\
x=-8-3t,\\
y=2+2t,\\
z=-7t\\
a=-3\\
b=2\\
c=-7\\
a(x-x_o)+b(y-y_o)+c(z-z_o)=0\\
-3(x-3)+2(y-(-2))-7(z-4)=0\\
3x-2y+7z=41"